# RudolfV: A Foundation Model by Pathologists for Pathologists

### 23 Jan 2024 v2

@@ Abstract
Histopathology plays a central role in clinical medicine and biomedical research. While artificial intelligence shows promising results on many pathological tasks, generalization and dealing with rare diseases, where training data is scarce, remains a challenge. Distilling knowledge from unlabeled data into a foundation model before learning from, potentially limited, labeled data provides a viable path to address these challenges. In this work, we extend the state of the art of foundation models for digital pathology whole slide images by semi-automated data curation and incorporating pathologist domain knowledge.
Specifically, we combine computational and pathologist domain knowledge (1) to curate a diverse dataset of 103k slides corresponding to 750 million image patches covering data from different fixation, staining, and scanning protocols as well as data from different indications and labs across the EU and US, (2) for grouping semantically similar slides and tissue patches, and (3) to augment the input images during training. We evaluate the resulting model on a set of public and internal benchmarks and show that although our foundation model is trained with an order of magnitude less slides, it performs on par or better than competing models. We expect that scaling our approach to more data and larger models will further increase its performance and capacity to deal with increasingly complex real world tasks in diagnostics and biomedical research.
@@ 1 Introduction
Despite artificial intelligence (AI) having led to promising proof-of-concepts and applications (e.g., [1, 2, 3, 4, 5]), generalization and robustness are still a challenge when translating applications into clinical setups. Moreover, limited numbers of samples available for training pose a particular challenge in medicine, because the distribution of disease frequencies is highly skewed with only relatively few frequent and a large number of rare diseases. Therefore, generating enough labeled data to sufficiently cover observed variations regarding morphology, staining, scanner type, lab, and biology is (prohibitively) expensive.
Figure 1: (1) Curated data: We curated a dataset of 103k slides with diverse properties. From these slides we extracted 750 million image patches. (2) Combining computational and pathologist expertise: With Pathologist input, we grouped the slides into $26$ groups, and - starting from a computer vision based clustering - the image patches into semantically meaningful tissue clusters. (3) AI training: We adapted DINOv2 training to sample a specific distribution of slide groups and tissue clusters and extend the augmentations with stain variations. (4) Applications: The resulting foundation model can be used for various applications in digital pathology.
Thus, while learning from labeled data will remain the standard for most AI use cases, strategies will be required to address the above outlined challenges. Here, foundation models [6] may contribute to a solution. This technique refers to pre-training a so-called foundation model on unlabeled data and to subsequently adapting it to learn from, potentially limited, labeled data. It typically uses the self-supervised learning (SSL, e.g., [7, 8, 9, 10]) paradigm and has shown breakthrough results in other domains when employed with an abundance of unlabelled data coupled with very large models (e.g., [11, 12, 13, 14, 8]).
@@
Our work is based on the assumption that optimally curated and more diverse data in connection with domain knowledge can improve SSL training. We therefore integrate pathologist knowledge in a number of key aspects. Specifically, we use pathologist input (1) to curate a dataset of 103k slides that covers diverse histological samples from different fixation, staining, and scanning protocols as well as data from different indications and labs from the EU and US, (2) for grouping semantically similar slides and tissue patches as well as (3) for augmenting the input images during training. Additional technical contributions are outlined in the methods section. We evaluate the resulting model on a set of public and internal benchmarks and compare it to latest related work.
@@ 3.1 Data Curation
Figure 2: Distribution of tissue origin, where available, in terms of organ systems in our dataset of 103,849 whole slide images (WSI).
  • Cases and slides: The dataset consists of 103,849 slides comprising 35,784 cases. 73,774 (71%) slides have been chosen from our archive and 30,075 (29%) slides are from The Cancer Genome Atlas (TCGA).
@@
  • Staining: To the best of our knowledge, we are the first to combine H&E and non-H&E stainings. Our dataset contains 3 broad staining categories: H&E (70%), IHC (10%), and other (20%). Subdividing each staining category further results in 97 unique staining types. Examples are H&E, ER, PR, PD-L1, HER-2, Giemsa, PAS, Gomori, and others.
@@
From each slide we extract patches within given tissue boundaries derived from an in-house model. The patch size is $256\times 256$ pixels at $0.5$ mpp. This results in $791$ million image patches in total for further processing and training.
@@ 3.2 Slide Grouping and Tissue Clustering
Histological slides typically come with metadata available for data curation. Based on expert input as well as the sample’s lab of origin, tissue type, and staining, all slides were assigned to one of 26 groups. These groups are selected to maximize homogeneity within the groups and heterogeneity across the groups. During training we sample patches uniformly from these slide groups.
In contrast, individual tissue patches typically have no tissue specific metadata available. For optimizing patch sampling during training, we therefore create clusters based on computer vision image features and pathologist expertise. First, we extract 36 standard computer vision image features for all 791 million image patches. These include image statistics such as mean and standard deviation for individual channels in common color modes such as RGB, LAB, HSV, and - tailored towards digital pathology - HED [35]. In contrast to [8], we do not use deep network features to reduce the computational burden. Next, we further reduce computational overhead by sub-sampling 500 patches for each slide and perform k-means clustering ($k=100$) on the respective image features. We propagate the labels of the clustering via K-NN to assign a cluster label to each of the 791 million patches. Finally, board-certified pathologists filter and aggregate the 100 image clusters assigning a description and a sampling probability (importance weight) to each cluster. This resulted in 9 distinct human-interpretable tissue clusters and 751 million image patches after filtering irrelevant clusters. During training, we sample patches according to the sampling probability of the clusters. Within the clusters, we sample equally from the 26 slide groups.
@@ 3.4 DINOv2
We build upon the official DINOv2 implementation [8] with registers [38]. DINOv2 learns representations via self-distillation. As [31] has shown that self-supervised histopathology models can benefit from pre-training on natural images, we initialize the student model with the distilled DINOv2 version pre-trained on LVD-142M [8]. This has shown to improve performance in smaller experimental setups. We train a ViT-L/14 with a batch size of 960 on 16 A100-40GB GPUs for 625k iterations.
@@ 4.1 Tasks
The benchmarks contain 8 different patch-based downstream tasks covering different tissue types, prediction tasks, and staining protocols.
@@
HE (8 classes)
The task is to predict the cell type at the center of an image patch ($128\times 128$ pixels at 0.5mpp). The eight classes are carcinoma cell, epithelial cell, plasma cell, fibroblast, lymphocyte, granulocyte, macrophage, and other. This internal benchmark dataset consists of 84,017, 6,695, and 14,532 board-certified pathologist annotated image patches taken from a larger corpora, split into train, validation, and test set respectively. The data is from 361 NSCLC slides of different tissue locations and labs.
PCAM, MHIST, CRC-100K benchmarks are an out-of-distribution validation for our model as the data was generated by separate labs. The internal HE and IHC tasks are an out-of-sample validation for our model as we carefully separated cases from the pre-training and task datasets. For TIL Det., MSI CRC, MSI STAD there is a potential overlap between pre-training and task data, but we assume the impact to be negligible. Please note, similarly for CTransPath [15] and Phikon [16] it is unclear if pre-training and TIL Det., MSI CRC, MSI STAD task data overlap.
@@ 4.2 Results
We evaluated all public benchmarks with linear probing and internal benchmarks with linear probing and finetuning. In line with related work [18], we do not use any augmentations on public benchmarks, while on the internal benchmarks, we use color augmentations during training. During linear probing, we resize the images to 224px and fit a linear classifier with l2-regularization on the CLS token [8] representations. The l2-regularization parameter is selected based on a hold-out validation set (10% of the official training set or the official validation set, if existing). During finetuning we extend the CLS token with another linear layer and use mini-batch stochastic gradient descent with a logistics loss.
@@
ResNet50 ImageNet 0.818 0.783 0.838 0.915 0.653 0.664 0.813 0.779
ViT-L/16 ImageNet 0.852 0.792 0.843 0.924 0.669 0.671 0.829 0.792
DINO ViT-S/8 [30] 0.901 * 0.779 * 0.896 0.943 0.745 0.756 0.859 0.837
CTransPath [15] 0.858 * 0.789 * 0.879 * 0.930 0.694 0.726 0.842 0.813
Phikon [16] 0.909 * 0.793 * 0.895 * 0.946 0.733 0.751 0.866 0.838
Virchow [18] 0.927 * 0.827 * 0.948 * - - - 0.901 -
Ours: RudolfV 0.928 0.827 0.960 0.939 0.766 0.768 0.905 0.865
The results of the public benchmarks in Table 1 show that our proposed approach yields better performance on all tasks except on the TIL detection task, where Phikon [16] yields best performance. The comparison covers all pathology SSL models that are publicly available at the time of publication. Virchow [18] is not publicly available and could therefore not be included in all benchmarks  [^1] We compare to the results in version 4 of the Virchow preprint [18]. While version 5 does not indicate a change in approach or model, it changes the evaluation procedure of all tasks and is thus not directly comparable anymore. Changes can be observed here: https://arxiv.org/abs/2309.07778..
@@
Table 2: Evaluation on internal cell classification benchmarks. We report balanced accuracy.
IHC (4 classes) HE (8 classes)
Model Finetuning Linear Finetuning Linear
ResNet50 ImageNet 0.6507 0.2914 0.6112 0.2143
ViT-L/16 ImageNet 0.8237 0.4234 0.7352 0.3305
Phikon [16] 0.8357 0.6373 0.8361 0.607
Ours: RudolfV 0.8745 0.6567 0.8864 0.7352
Table 2 shows evaluation results on internal benchmarks. Our proposed approach yields better results, both, for linear probing and for finetuning. For linear probing the model is up to 13 p.p. and for finetuning it is up to 5 p.p. better than the second best model.

### 12 Mar 2024 v3

@@ Abstract
Histopathology plays a central role in clinical medicine and biomedical research. While artificial intelligence shows promising results on many pathological tasks, generalization and dealing with rare diseases, where training data is scarce, remains a challenge. Distilling knowledge from unlabelled data into a foundation model before learning from, potentially limited, labelled data provides a viable path to address these challenges. In this work, we extend the state of the art of foundation models for digital pathology whole slide images by semi-automated data curation and incorporating pathologist domain knowledge.
Specifically, we combine computational and pathologist domain knowledge (1) to curate a diverse dataset of 133k slides corresponding to 1.2 billion image patches covering data from different fixation, staining, and scanning protocols as well as data from different indications and labs across the EU and US, (2) for grouping semantically similar slides and tissue patches, and (3) to augment the input images during training. We evaluate the resulting model on a set of public and internal benchmarks and show that although our foundation model is trained with an order of magnitude less slides, it performs on par or better than competing models. We expect that scaling our approach to more data and larger models will further increase its performance and capacity to deal with increasingly complex real world tasks in diagnostics and biomedical research.
@@ 1 Introduction
Despite artificial intelligence (AI) having led to promising proof-of-concepts and applications (e.g., [1, 2, 3, 4, 5]), generalization and robustness are still a challenge when translating applications into clinical setups. Moreover, limited numbers of samples available for training pose a particular challenge in medicine, because the distribution of disease frequencies is highly skewed with only relatively few frequent and a large number of rare diseases. Therefore, generating enough labelled data to sufficiently cover observed variations regarding morphology, staining, scanner type, lab, and biology is (prohibitively) expensive.
Figure 1: (1) Curated data: We curated a dataset of 133k slides with diverse properties. From these slides we extracted 1.2 billion image patches. (2) Combining computational and pathologist expertise: With Pathologist input, we grouped the slides into $31$ groups, and - starting from a computer vision based clustering - the image patches into semantically meaningful tissue clusters. (3) AI training: We adapted DINOv2 training to sample a specific distribution of slide groups and tissue clusters and extend the augmentations with stain variations. (4) Applications: The resulting foundation model can be used for various applications in digital pathology.
Thus, while learning from labelled data will remain the standard for most AI use cases, strategies will be required to address the above outlined challenges. Here, foundation models [6] may contribute to a solution. This technique refers to pre-training a so-called foundation model on unlabelled data and to subsequently adapting it to learn from, potentially limited, labelled data. It typically uses the self-supervised learning (SSL, e.g., [7, 8, 9, 10]) paradigm and has shown breakthrough results in other domains when employed with an abundance of unlabelled data coupled with very large models (e.g., [11, 12, 13, 14, 8]).
@@
Our work is based on the assumption that optimally curated and more diverse data in connection with domain knowledge can improve SSL training. We therefore integrate pathologist knowledge in a number of key aspects. Specifically, we use pathologist input (1) to curate a dataset of 133k slides that covers diverse histological samples from different fixation, staining, and scanning protocols as well as data from different indications and labs from the EU and US, (2) for grouping semantically similar slides and tissue patches as well as (3) for augmenting the input images during training. Additional technical contributions are outlined in the methods section. We evaluate the resulting model on a set of public and internal benchmarks and compare it to latest related work.
@@ 3.1 Data Curation
Figure 2: Distribution of tissue origin, where available, in terms of organ systems in our dataset of 133k whole slide images (WSI).
  • Cases and slides: The dataset consists of 133,998 slides comprising 34,103 cases. 108,433 (81%) slides have been chosen from our archive and 26,565 (19%) slides are from The Cancer Genome Atlas (TCGA).
@@
  • Staining: To the best of our knowledge, we are the first to combine H&E and non-H&E stainings. Our dataset contains 3 broad staining categories: H&E (68%), IHC (15%), and other (17%). Subdividing each staining category further results in 129 unique staining types. Examples are H&E, ER, PR, PD-L1, HER-2, Giemsa, PAS, Gomori, and others.
@@
From each slide we extract patches within given tissue boundaries derived from an in-house model. The patch size is $256\times 256$ pixels at $0.5$ mpp. This results in $1.25$ billion image patches in total for further processing and training.
@@ 3.2 Slide Grouping and Tissue Clustering
Histological slides typically come with metadata available for data curation. Based on expert input as well as the sample’s lab of origin, tissue type, and staining, all slides were assigned to one of 31 groups. These groups are selected to maximize homogeneity within the groups and heterogeneity across the groups. During training we sample patches uniformly from these slide groups.
In contrast, individual tissue patches typically have no tissue specific metadata available. For optimizing patch sampling during training, we therefore create clusters based on computer vision image features and pathologist expertise. First, we extract 36 standard computer vision image features for all 1.25 billion image patches. These include image statistics such as mean and standard deviation for individual channels in common color modes such as RGB, LAB, HSV, and - tailored towards digital pathology - HED [35]. In contrast to [8], we do not use deep network features to reduce the computational burden. Next, we further reduce computational overhead by sub-sampling 500 patches for each slide and perform k-means clustering ($k=100$) on the respective image features. We propagate the labels of the clustering via K-NN to assign a cluster label to each of the 1.25 billion patches. Finally, board-certified pathologists filter and aggregate the 100 image clusters assigning a description and a sampling probability (importance weight) to each cluster. This resulted in 9 distinct human-interpretable tissue clusters and 1.2 billion image patches after filtering irrelevant clusters. During training, we sample patches according to the sampling probability of the clusters. Within the clusters, we sample equally from the 31 slide groups.
@@ 3.4 DINOv2
We build upon the official DINOv2 implementation [8] with registers [38]. DINOv2 learns representations via self-distillation and the final embedding representation is defined as in [18]. As [31] has shown that self-supervised histopathology models can benefit from pre-training on natural images, we initialize the student model with the distilled DINOv2 version pre-trained on LVD-142M [8]. This has shown to improve performance in smaller experimental setups. We train a ViT-L/14 with a batch size of 960 for 625k iterations.
@@ 4.1 Tasks
The benchmarks contain 9 different patch-based downstream tasks covering different tissue types, prediction tasks, and staining protocols.
PCAM, MHIST, CRC-100K benchmarks are an out-of-distribution validation for our model as the data was generated by separate labs. The internal HE and IHC tasks are an out-of-sample (or out-of-distribution, if indicated) validation for our model as we carefully separated cases from the pre-training and task datasets. For TIL Det., MSI CRC, MSI STAD there is a potential overlap between pre-training and task data, but we assume the impact to be negligible. Please note, similarly for CTransPath [15] and Phikon [16] it is unclear if pre-training and TIL Det., MSI CRC, MSI STAD task data overlap.
@@
Pan-indication HE cell classification (5 indications, 8 classes)
The task is to predict the cell type at the center of an image patch ($128\times 128$ pixels at 0.5mpp). The indications are urothelial cancer (UCa), breast cancer (BrCa), colorectal cancer (CRC), hepatocellular carcinoma (HCC) and non-small-cell lung cancer (NSCLC). The eight prediction classes are carcinoma cell, epithelial cell, plasma cell, fibroblast, lymphocyte, granulocyte, macrophage, and other. This internal benchmark dataset consists of 74,034, 24,692, and 70,198 board-certified pathologist annotated image patches taken from a larger corpora, split into train, validation, and test set respectively. The annotations are gathered from 347 slides of different tissue locations and labs. We created a hold out set for each indication consisting of cases not seen during training. For out-of-distribution evaluation we created a additional hold out dataset based on NSCLC cases from a different cohort and lab.
Pan-indication HE segmentation (3 indications, 5 classes)
The task is semantic segmentation on an internal cohort, trained and evaluated on three indications: colorectal cancer (CRC), hepatocellular carcinoma (HCC) and non-small-cell lung cancer (NSCLC). From each slide we extract annotated patches at 0.5mpp. Depending on the model the patch input size was different: 224 pixels for the ImageNet based ViT and 336 pixels for the model from [16] and our model. For every pixel on the patch the model predicts one of the following classes: Carcinoma, Stroma, Ephitelial tissue, Necrosis or other. The training dataset originates from 456 slides of different tissue locations and labs, and is split into 30,553 train and 7,895 dev patches. We created a hold out set with a total of 11,800 patches, consisting of cases not seen during training. For out-of-distribution evaluation we created an additional hold out dataset based on NSCLC cases from a different cohort and lab.
@@ 4.2 Results
Public Classification Benchmarks: We evaluated all public benchmarks with linear probing. In line with related work [18], we do not use any augmentations on public benchmarks. During linear probing, we resize the images to 224 pixels and fit a linear classifier on the representations. We use the Adam optimizer with learning rate $55\times 10^{-}$ 5, batch size 128 and cosine learning rate schedule.
@@
ResNet50 ImageNet 0.833 0.806 0.849 0.915 0.653 0.664 0.829 0.786
ViT-L/16 ImageNet 0.852 0.796 0.847 0.924 0.669 0.671 0.832 0.793
Kang et al. [30] 0.918 * 0.771 * 0.949 * 0.943 0.745 0.756 0.879 0.847
CTransPath [15] 0.872 * 0.817 * 0.840 * 0.930 0.694 0.726 0.843 0.813
Phikon [16] 0.906 * 0.795 * 0.883 * 0.946 0.733 0.751 0.861 0.836
Virchow [18] [^1] We compare to the latest results of the Virchow preprint ([18], version 5). 0.933 * 0.834 * 0.968 * - - - 0.912 -
Ours: RudolfV 0.944 0.821 0.973 0.943 0.755 0.788 0.913 0.871
The results of the public benchmarks in Table 1 show that our proposed approach yields better performance on all tasks except on the TIL detection and MHIST task, where Phikon [16] and Virchow [18] yield best performance respectively. Virchow [18] is not publicly available and could therefore not be included in all benchmarks.
@@
Internal Classification Benchmarks: We evaluated all internal benchmarks with linear probing and finetuning. We extend the model embeddings with an additional linear layer and use mini-batch stochastic gradient descent with a logistic loss to train the models. For linear probing, the encoder weights are not updated and we use color augmentations during training for the HE task.
Table 2: Model finetuning: Evaluation on internal cell classification benchmarks. All models are fully finetuned and we report macro-averaged balanced accuracy. OOD denotes an out-of-distribution evaluation.
HE (5 indications, 8 classes) IHC
Model UCa BrCa CRC HCC NSCLC NSCLC OOD Average (4 classes)
ResNet50 ImageNet 0.741 0.637 0.655 0.675 0.675 0.638 0.670 0.651
ViT-L/16 ImageNet 0.759 0.659 0.705 0.711 0.711 0.700 0.701 0.824
Phikon [16] 0.867 0.771 0.833 0.843 0.870 0.883 0.845 0.836
Ours: RudolfV 0.904 0.888 0.867 0.890 0.923 0.931 0.901 0.870
Table 3: Linear probing: Evaluation on internal cell classification benchmarks. For each model the encoder is frozen and only the linear classification head is trained. We report macro-averaged balanced accuracy. OOD denotes an out-of-distribution evaluation.
HE (5 indications, 8 classes) IHC
Model UCa BrCa CRC HCC NSCLC NSCLC OOD Average (4 classes)
ResNet50 ImageNet 0.282 0.239 0.217 0.222 0.232 0.225 0.236 0.291
ViT-L/16 ImageNet 0.470 0.402 0.382 0.371 0.423 0.316 0.394 0.423
Phikon [16] 0.776 0.648 0.679 0.582 0.744 0.682 0.685 0.637
Ours: RudolfV 0.848 0.689 0.754 0.644 0.800 0.794 0.755 0.688
Table 2 and 3 show the evaluation results on internal HE classification benchmarks. Our proposed approach yields better results for all evaluation sets, both, for linear probing and for finetuning. For linear probing the model is on average 7 p.p. and for finetuning it is on average 5.6 p.p. better than the second best model Phikon[16]. On the out-of-distribution cohort the difference changes to 11.2 p.p and 4.8 p.p, respectively.
Table 2 and 3 also show the evaluation results on internal IHC classification benchmarks. Our proposed approach yields better results, both, for linear probing and for finetuning. For linear probing the model is 5.1 p.p. and for finetuning it is 3.4 p.p. better than the second best model Phikon[16].
Internal Segmentation Benchmarks: We evaluated all internal benchmarks in two modes, once with a frozen feature extractor and once with finetuning the entire model. We use a multi-layer segmentation head that operates exclusively on the patch embeddings of the encoder. We first bottleneck the embedding dimensionality, similar to [48], and then increase the spatial resolution by a factor of half the mini-patch size, using a bi-linear layer. Finally, we add a convolutional layer with the correct number of output channels, and upsample by a factor of two.
Table 4 and 5 show the evaluation results on internal segmentation benchmarks. Our proposed approach yields better results for all evaluation sets, both, for a frozen encoder and for finetuning. With a frozen encoder the model is on average 2.5 p.p. and when finetuning the whole model it is u 2.6 p.p. better than the second best model Phikon[16]. On the out-of-distribution cohort the difference increases to 5.1 p.p and 6.5 p.p, respectively.
Table 4: Model finetuning: Evaluation on internal segmentation benchmarks. Every model is fully finetuned. We report macro-averaged balanced accuracy. OOD denotes an out-of-distribution evaluation.
HE (3 indications, 5 classes)
Model NSCLC NSCLC OOD CRC HCC Average
ViT-L/16 ImageNet 0.786 0.669 0.888 0.812 0.789
Phikon [16] 0.905 0.778 0.945 0.897 0.881
Ours: RudolfV 0.912 0.843 0.962 0.910 0.907
Table 5: Frozen encoder: Evaluation on internal segmentation benchmarks. For each model the encoder is frozen and only the segmentation head is trained. We report macro-averaged balanced accuracy. OOD denotes an out-of-distribution evaluation.
HE (3 indications, 5 classes)
Model NSCLC NSCLC OOD CRC HCC Average
ViT-L/16 ImageNet 0.703 0.577 0.806 0.723 0.702
Phikon [16] 0.891 0.782 0.931 0.891 0.874
Ours: RudolfV 0.910 0.833 0.946 0.908 0.899